ndard Chartered Bank Zambia PLC (Stanchart) has 100,000 shares in issue at the beginning of 2020. On 30th June 2020, the bank issues 50,000 more shares at full market price of K2.3. The nominal price per share of the initial 100,000 shares was K1. The 2019 earnings before tax were K250,000 (2019: K330,000), with an effective tax rate of 30%.REQUIRED:Identify an appropriate company on the LUSE whose performance will be compared to that of Stanchart. Now analyse Stanchart’s performance for 2020 with reference to the following question.Calculate the earnings per share (EPS) as at 31st December 2019 and 2020 correct to two decimal places for Stanchart.[10 Marks]Calculate the earnings per share (EPS) as at 31st December 2019 and 2020 correct to two decimal places for the comparative company.[10
1
Expert's answer
2021-09-04T15:18:14-0400

Solution:

An appropriate company on the LUSE whose performance will be compared to that of Stanchart is Zambia National Commercial Bank plc.

 

a.). EPS = "\frac{Net\; Income}{Outstanding\; Shares}"

EPS 2019:

Net income 2019 = 330,000

Income tax 2019 = 30"\% \times" 330,000 = 99,000

Net Income after tax = 330,000 – 99,000 = 231,000

Total outstanding shares 2019 = 100,000 shares


EPS 2019 = "\frac{231,000}{100,000} = 2.31"


EPS 2019 = K2.31

 

EPS 2020:

Net income 2020 = 250,000

Income tax 2019 = 30"\% \times" 250,000 = 75,000

Net Income after tax = 250,000 – 75,000 = 175,000

Total outstanding shares = 100000 + 50000 = 150,000 shares


EPS 2019 = "\frac{175,000}{150,000} = 1.17"


EPS 2020 = K1.17

 

b.). EPS for the comparative company: Zambia National Commercial Bank plc.

EPS 2019:

Net income 2019 after tax = 200,056,000

Total outstanding shares 2019 = 1,443,750,000 shares


EPS 2019 = "\frac{200,056,000}{1,443,750,000} = 0.14"


EPS 2019 = 0.14

 

EPS 2020:

Net income 2020 after tax = 206,658,000

Total outstanding shares 2020 = 1,443,750,000 shares


EPS 2020 = "\frac{206,658,000}{1,443,750,000} = 0.14"


EPS 2020 = K 0.14